Assume the nominal rate of return is 9.78% and the inflation rate is 2.82%. Find the real rate of return using the exact formula.
Answer #1

Nominal Rate of Return (NRR) = 9.78 % and Inflation Rate (I) = 2.82 %

The real rate of return (RRR) and NRR are related as per the formula given below:

RRR + 1 = (1+NRR) / (1+Infation Rate) = (1.0978) / (1.0282) ~ 1.0677

RRR = 0.0677 or 6.77 %